OpenStudy (anonymous):

How many solutions does the equation have? a+5=1/5(15+5a) A. one solution B. infinite solutions C. no solution

OpenStudy (anonymous):

Why don't you open this first

OpenStudy (anonymous):

what @KissMyAxe?

jhonyy9 (jhonyy9):

so this mean that there are 15+5a a+5 = -------- so then factoriz out the 5 from 15 and 5a and will get 5 5(3+a) a+5 = -------- simplifie the 5 and get 5 a+5= 3+a subtract from both sides a and will get a-a+5=3+a-a 5=3 so do you know what mean this result ?

OpenStudy (anonymous):

the answer is A?

OpenStudy (anonymous):

?????

jhonyy9 (jhonyy9):

why A one solution when the unknowed a. will be eliminated ?????

jhonyy9 (jhonyy9):

so i think C.

OpenStudy (anonymous):

ooo ok thanks :D

jhonyy9 (jhonyy9):

np was my pleasure good luck bye